OpenStudy (anonymous):

In a given day, MyTCo, Inc., sold 119 T-shirts. Plain shirts sold for $15.50 each and striped ones for $18.50 each. The company took in $1988.50 . How many of the plain shirts were sold? 48 65 71 87

rebeccaxhawaii (rebeccaxhawaii):

t+p=119 15.50p+18.50t=1988.50

OpenStudy (anonymous):

So you can write this as the equation of\[15.5x+18.5y=1988.5\]and\[x+y=119\]You can solve this as a system of equations, I recommend substitution

rebeccaxhawaii (rebeccaxhawaii):

yes substitution would be easier

OpenStudy (anonymous):

what do i substitute?

OpenStudy (anonymous):

@rebeccaxhawaii

rebeccaxhawaii (rebeccaxhawaii):

y=-x+119 replace it into the equation for y 15.5x+18.5y=1988.5 15.5x+18.5(-x+119)=1988.5

OpenStudy (anonymous):

I came out with 71 I am not sure if that is right but thank you so much for your help @rebeccaxhawaii